OpenStudy (anonymous):

Miguel and his wife are each starting a saving plan. Miguel will initially set aside $75 and then add $30.65 every week to the savings. The amount A (in dollars) saved this way is given by the function A= 75+30.65N, where N is the number of weeks he has been saving. His wife will not set an initial amount aside but will add $55.85 to the savings every week. The amount B (in dollars) saved using this plan is given by the function B = 55.85N Let T be total amount (in dollars) saved using both plans combined. Write an equation relating to N. Simplify you answer as much as possible.

OpenStudy (anonymous):

T = N ((75 + 30.65) + (55.85))

OpenStudy (anonymous):

I wasn't quite sure where to start so I'm not sure if i'm on the right track or not =o @amistre64

OpenStudy (anonymous):

T= N (105.65) + (55.85) T= N (161.5)

OpenStudy (anonymous):

OR is it T = N((30.65) + (55.85)) + 75 ( since 75$ deposit was only made once) T = 86.05N + 75

OpenStudy (amistre64):

A + B you cant factor our an N from 75 since it has no N to start with

OpenStudy (amistre64):

75 + 0 is the initial amount thats not changeing 31 + 56 is added each N, so this amount changes the value in the account over time.